# SONAR: A Synthetic AI-Audio Detection Framework and Benchmark

### v3

@@ 5 Discussion
Data license considerations.. Since our dataset is sourced from various models, each may be subject to different distribution licenses and usage restrictions. Throughout the data collection process, we ensured strict adherence to all relevant usage policies. We have also made the dataset accessible to everyone, either directly via the provided link or indirectly through the original sources. However, as these policies are frequently refined and updated, we will ensure that our published dataset remains in full compliance with the latest regulations. Additionally, we will reference the usage policies of the respective API providers to ensure that users are informed of any potential restrictions.
6 Conclusion
In this paper, we presented SONAR, a framework providing a comprehensive evaluation for distinguishing state-of-the-art AI-synthesized auditory content. SONAR introduces a novel evaluation dataset sourced from 9 diverse audio synthesis platforms, including leading TTS service providers and state-of-the-art TTS models. To the best of our knowledge, SONAR is the first platform that provides uniform, comprehensive, informative, and extensible evaluation of deepfake audio detection models. Leveraging SONAR, we conducted extensive experiments to analyze the generalizability limitations of current detection methods. We found that foundation models demonstrate stronger generalization capabilities, given their massive model size scale and pertaining data. We further explored the potential of few-shot fine-tuning to improve generalization and demonstrated its efficiency and effectiveness. We envision that SONAR will serve as a valuable benchmark to facilitate research in AI-audio detection and highlight directions for further improvement.
Acknowledgments and Disclosure of Funding
@@ 6 Conclusion
Appendix A Appendix
A.1 Broad Impacts
Positive Impacts. On the positive side, AI-synthesized audio/speech has the potential to revolutionize content creation in various sectors, including education, entertainment, and accessibility. In education, AI-synthesized audios and speeches enables production of customized content that meets diverse learning needs and languages, improving access and inclusivity. For entertainment, they can offer novel experiences by generating dynamic dialogues in games or virtual reality, enriching user engagement and creativity.
Furthermore, AI-synthesized audios and speeches also enhance accessibility by producing speech in various languages or dialects, bridging communication gaps and making information more accessible to non-native speakers or those with liabilities. Additionally, the technology can help preserve lesser-spoken languages and dialects at risk of extinction by creating archives of AI-generated speeches and narratives.
In conclusion, while AI-synthesized audios and speeches offer exciting opportunities for content creation and accessibility, it is essential to address the ethical and societal challenges associated with its use. Collaborative efforts among researchers, developers, and policymakers are crucial to leveraging AI-synthesized audio and speech benefits responsibly while mitigating its risks, ensuring the technology serves to enhance human communication and creativity positively and responsibly.
A.2 Implementation Details
Table 7: Hyperparemeters
config value
optimizer Adam
optimizer momentum $\beta_{1}=0.9$, $\beta_{2}=0.999$
weight decay 1e-4
epochs 40
warmup epochs 0
scheduler cosine decay
Table 7 presents the hyperparameters for training AASIST, RawNet2, RawGAT-ST, LCNN, and Spec.+ResNet. We train AASIST, RawNet2, and RawGAT-ST with a learning rate of 0.0001 and LCNN and Spec.+ResNet with a learning rate of 0.0003. The batch size for AASIST, RawNet2, RawGAT-ST, LCNN, and Spec.+ResNet are 64, 256, 32, 512, and 256, respectively. All input audios are resampled to a 16kHz sampling rate and converted into raw waveforms consisting of 64,000 samples (approximately 4 seconds). Audios longer than 4 seconds are randomly trimmed, while those shorter than 4 seconds are repeated and padded to meet the 4-second duration.
For few-shot fine-tuning, models are fine-tuned for 30 epochs with a learning rate of 0.00001 and a weight decay of 0.00005.

### v4

@@ 5. Discussion
Data license considerations.. Since our dataset is sourced from various models, each may be subject to distinct distribution licenses and usage restrictions. Throughout the data collection process, we strictly adhered to all relevant usage policies. The dataset is made accessible either directly through the provided link or indirectly via the original sources. To account for the evolving nature of these policies, we are committed to keeping the published dataset fully compliant with the latest regulations. Additionally, we will reference the usage policies of the respective API providers to inform users about any potential restrictions.
6. Conclusion
In this paper, we presented SONAR, a framework providing a comprehensive evaluation for distinguishing state-of-the-art AI-synthesized auditory content. SONAR introduces a novel evaluation dataset sourced from 9 diverse audio synthesis platforms, including leading TTS service providers and state-of-the-art TTS models. To the best of our knowledge, SONAR is the first platform that provides uniform, comprehensive, informative, and extensible evaluation of deepfake audio detection models. Leveraging SONAR, we conducted extensive experiments to analyze the generalizability limitations of current detection methods. We found that speech foundation models demonstrate stronger generalization capabilities across datasets and languages, given their massive model size scale and pertaining data. We also suggest that the primary challenges in audio deepfake detection are more closely tied to the realism and quality of synthetic audio rather than language-specific characteristics. In addition, we further explored the potential of few-shot fine-tuning to improve generalization and demonstrated its efficiency and effectiveness. We envision that SONAR will serve as a valuable benchmark to facilitate research in AI-audio detection and highlight directions for further improvement.
Acknowledgments